[ the perimeter of a rectangle is 120 feet. if the length is four times the width, what are the dimensions of the rectangle? ]
Answer by TimothyLamb(4379)

---
x = length
y = width
---
2x + 2y = 120
x = 4y
---
put the system of linear equations into standard form
---
2x + 2y = 120
x - 4y = 0
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 48 ft
y = width = 12 ft
---
system matrix
2 2
1 -4
---
inverse of system matrix
0.4 0.2
0.1 -0.2
---
determinant of system matrix = -10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php